ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

12. Probability and Statistics

1991 MS Classification: 60-01, 60A05 Vinogradova, Natalya, Karelian State Pedagogical University, Petrozavodsk, Russia Classical Definition of Probability Basic questions of the theory of probability were included into Russian school programs only several years ago. It is necessary to find methods to teach the beginners. The classical definition of probability gives an opportunity to analyse the events‘ probabilities, operating with numbers. Problems with finite numbers of possible outcomes are very useful for the beginners, because they can help to realise the basic notions of probability. Reference: N. Vinogradova, The Theory of Probability for the beginners (in Russian), (1996).